OpenStudy (anonymous):

Solve for v s=(1/2)gt^2+vt

OpenStudy (isaiah.feynman):

First think of a way to get rid of that 1/2

OpenStudy (anonymous):

multiply 2 to each side.

OpenStudy (anonymous):

Unneeded, just subtract the leading term from both sides. Thus, you have: s - (1/2)gt^2 = vt Now what can you do to isolate v on the right?

OpenStudy (isaiah.feynman):

Now it becomes 2s=gt^2+vt.

OpenStudy (isaiah.feynman):

@bloopman you should have said "This is an alternative way" rather than unneeded. Both methods still give the same answer.

OpenStudy (anonymous):

But of course. I apologize.

OpenStudy (anonymous):

do i subtract vt from each side? getting 2s-vt=gt^2?

OpenStudy (isaiah.feynman):

Nope subtract gt^2 from both sides.

OpenStudy (isaiah.feynman):

2s-gt^2=vt, v= 2s-gt^2/t

OpenStudy (anonymous):

it says the answer is v=2s-gt^2/2t?

OpenStudy (isaiah.feynman):

Yes its correct. I forgot something, I saw my mistake. Your answer id right indeed!

OpenStudy (anonymous):

I don't understand where the 2t in the denom comes from though?

OpenStudy (anonymous):

Shifting the 1/2 down.

OpenStudy (isaiah.feynman):

Let me do the full correct calculation on paint. :)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Thank you!